# Syntomic cohomology of truncated Brown–Peterson spectra

### v1

@@ 6. Algebraic K-theory
In this section, we compute algebraic K-theory of $\operatorname{BP}\langle n\rangle$ at sufficiently low heights. We then prove Lichtenbaum—Quillen properties for algebraic K-theory of $\operatorname{BP}\langle n\rangle$. Finally, we prove that the telescope conjecture is true for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
@@ 6.1. Algebraic K-theory at low heights
$\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}[v_{3}]\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
as $\mathbb{F}_{p}[v_{3}^{d}]$-modules for some integer $d\geq 1$. Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
@@
as $\mathbb{F}_{p}[v_{3}^{\pm d}]$-modules for some integer $d\geq 1$. Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
@@ 6.2. Lichtenbaum–Quillen
For any integers $i_{0},\cdots,i_{n+1}$ such that the generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$ exists, the graded abelian group
$\pi_{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p,\cdots,v_{n+1})$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
is finite for any integers $i_{0},\cdots,i_{n+1}$. Choosing integers $i_{0},\cdots,i_{n+1}$ such that a generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$ exists, this immediately implies that
$\pi_{*}\operatorname{TC}(R)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
@@
the algebraic K-theory of $R$ has fp-type $n+1$ since $\pi_{*}\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ is also finite. ∎
@@
This implies the higher height analogue of the Lichtenbaum–Qullen conjecture as proposed by Ausoni and Rognes in [AR08].
@@
This follows from Theorem 6.4 by [MR99, Theorem 8.2(2)] as observed in [HW22, Theorem 3.1.3]. ∎
@@ 6.3. Telescope
We now note that our results also imply the telescope conjecture for topological cyclic homology of $\operatorname{BP}\langle n\rangle$.

### v2

@@ 6. Algebraic K-theory
In this section, we compute algebraic K-theory of $\operatorname{BP}\langle 2\rangle$ at primes $p\geq 5$. We then prove the Lichtenbaum—Quillen property for algebraic K-theory of $\operatorname{BP}\langle n\rangle$. Finally, we prove that the telescope conjecture holds for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
@@ 6.1. Algebraic K-theory at low heights
$\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\partial,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}[v_{3}]\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
as $\mathbb{F}_{p}[v_{3}^{d}]\langle\partial,\lambda_{1},\lambda_{2},\lambda_{3}\rangle$-modules for some integer $d\geq 1$. Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
@@
as $\mathbb{F}_{p}[v_{3}^{\pm d}]\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle$-modules for some integer $d\geq 1$. Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
@@ 6.2. Lichtenbaum–Quillen
For any integers $i_{0},i_{1},\cdots,i_{n+1}$ such that the generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ exists, the graded abelian group
$\pi_{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
@@
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
is finite for any integers $i_{0},i_{1},\cdots,i_{n+1}$ using the finite $p,v_{1},\cdots,v_{n+1}$-Bockstein filtrations. Choosing integers $i_{0},i_{1},\cdots,i_{n+1}$ such that the Moore spectrum $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ exists, this immediately implies that
$\pi_{*}\operatorname{TC}(R)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
@@
and the algebraic K-theory of $R$ has fp-type $n+1$ since $\pi_{*}\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ is also finite. ∎
@@
This implies the higher height analogue of the Lichtenbaum–Qullen conjecture as proposed by Ausoni and Rognes in [AR08] as part of the family of conjectures known as the redshift conjectures.
@@
This follows from Theorem 6.4 by [MR99, Theorem 8.2(2)] as observed in [HW22, Theorem 3.1.3] together with an arithmetic fracture square argument. ∎
@@ 6.3. Telescope
Since it is now known that the telescope conjecture is false in general [BHL+23], it becomes an interesting question to ask for which spectra $X$ is the localization map
$L_{n+1}^{f}X\to L_{n+1}X$
is an equivalence. For example, it was conjectured by [MR99, Conjecture 3.9] that the telescope conjecture holds for $fp$-spectra. We show that our results imply the telescope conjecture for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$, giving some support for this conjecture in light of Theorem 6.4, which implies that the algebraic K-theory of $\operatorname{BP}\langle n\rangle$ is an $fp$-spectrum.